two numbers total 64 and have a difference of 26. use x and y as an unknown number.

Answer by piglet162431(90)

Let's make an equation
y=x+26
x + x +26=64
2x+26=64
2x=38
x= 19
Now, to figure out what's y, use the first equation.
y=19+26
y=45
the numbers are 19 and 26